### v1

# MAS-Bench: A Unified Benchmark for Shortcut-Augmented Hybrid Mobile GUI Agents

Abstract
To enhance the efficiency of GUI agents on various platforms like smartphones and computers, a hybrid paradigm that combines flexible GUI operations with efficient shortcuts (e.g., API, deep links) is emerging as a promising direction. However, a framework for systematically benchmarking these hybrid agents is still underexplored. To take the first step in bridging this gap, we introduce MAS-Bench, a benchmark that pioneers the evaluation of GUI-shortcut hybrid agents with a specific focus on the mobile domain. Beyond merely using predefined shortcuts, MAS-Bench assesses an agent’s capability to autonomously generate shortcuts by discovering and creating reusable, low-cost workflows. It features 139 complex tasks across 11 real-world applications, a knowledge base of 88 predefined shortcuts (APIs, deep-links, RPA scripts), and 7 evaluation metrics. The tasks are designed to be solvable via GUI-only operations, but can be significantly accelerated by intelligently embedding shortcuts. Experiments show that hybrid agents achieve significantly higher success rates and efficiency than their GUI-only counterparts. This result also demonstrates the effectiveness of our method for evaluating an agent’s shortcut generation capabilities. MAS-Bench fills a critical evaluation gap, providing a foundational platform for future advancements in creating more efficient and robust intelligent agents. The project page is available at https://pengxiang-zhao.github.io/MAS-Bench.
1 Introduction
The rise of Large Language Models (LLMs) (et al. 2024; Zhao et al. 2023; Minaee et al. 2024; Chang et al. 2024; Xu et al. 2024) is driving the development of Graphical User Interface (GUI) Agents, enabling them to operate diverse digital platforms such as computers, web browsers, and smartphones (Liu et al. 2025b; Tang et al. 2025; Liu et al. 2025a). Early research on mobile agents primarily focused on replicating human-like flexibility through GUI-only interaction (Lee et al. 2025; Gou et al. 2024; Cheng et al. 2024; You et al. 2024; Zhang et al. 2024; Lu et al. 2025). This approach grants agents the generality to operate on any application, but it often overlooks the significant efficiency advantages offered by more direct methods.
As GUI agents are increasingly applied to complex tasks, enhancing their efficiency has become a central research focus. In this context, hybrid paradigms have emerged as a promising solution, demonstrating effectiveness across a wide range of platforms (Wang et al. 2025b; Zheng et al. 2025; Zhang et al. 2025a, b; Jiang et al. 2025; Lee et al. 2023). This approach combines the speed and reliability of “shortcuts” such as Application Programming Interface (API) calls, deep links, and Robotic Process Automation (RPA) scripts (Bridle and McCreath 2006; Guerreiro, Gamboa, and Jorge 2008; Kennedy and Everett 2011; Tripathi 2018; Ling, Gao, and Wang 2020; Agostinelli et al. 2022), with the flexibility of GUI operations. As Fig. 1 illustrates, a hybrid agent can bypass tedious, multi-step GUI operations by invoking a single, efficient shortcut, drastically reducing operational complexity and time.
Figure 1: Workflow of GUI Only vs. GUI-Shortcut Hybrid Agent. Shortcuts improve agent execution efficiency by bypassing GUI operations.
However, despite these advancements, a framework for systematically evaluating and benchmarking hybrid agents is still lacking (Le et al. 2020; Roffarello, Purohit, and Purohit 2024; Zhang et al. 2025a, b). This gap leaves the full potential of GUI-Shortcut agents far from being thoroughly explored and assessed. While this is a general challenge, the mobile domain represents a critical and demanding environment for progress.
Therefore, we present MAS-Bench, the first benchmark explicitly designed to evaluate GUI-shortcut hybrid mobile agents. The tasks within MAS-Bench are solvable entirely through GUI interactions, but integrating shortcuts can improve efficiency. MAS-Bench evaluates agents’ ability to make optimal decisions between different operational modes to enhance performance. Our experiments on MAS-Bench show that hybrid agents achieve up to a 64.1% success rate, a significant improvement over the 44.6% achieved by their GUI-only counterparts, and execute tasks with over 40% greater efficiency.
Furthermore, we introduce a novel framework to evaluate agents’ capacity to generate new shortcuts from interaction. This framework integrates agent-generated shortcuts into a standard baseline agent and measures subsequent task performance. Our findings reveal a performance gap: while our predefined shortcuts prove highly reliable (100% success rate), agent-generated shortcuts lag in robustness and efficiency. Dynamic shortcuts demonstrate significant potential. Despite their task completion rate of 38% (baseline 43%), they offer the highest efficiency, highlighting them as a promising direction for future research into robust shortcut generation. Our contributions are threefold:
  • We introduce MAS-Bench, the first benchmark for systematically evaluating GUI-shortcut hybrid mobile agents. It comprises 139 complex tasks spanning 11 real-world applications, supported by a knowledge base of 88 predefined shortcuts and 7 distinct evaluation metrics.
  • We establish baselines with several agents on MAS-Bench, demonstrating that those leveraging GUI-shortcut hybrid operations achieve significantly higher success rates and efficiency than their GUI-only counterparts.
  • We propose the first framework to evaluate an agent’s ability to generate shortcuts from interaction. Our experiments reveal a substantial gap between agent-generated shortcuts and predefined ones in efficiency and robustness, highlighting a key area for future research.
2 Related Work
Mobile Task Automation.
Mobile task automation evolves from methods based on predefined scripts to more intelligent and adaptive agents driven by LLMs (Kirubakaran and Karthikeyani 2013; Amalfitano et al. 2014; Linares-Vásquez, Moran, and Poshyvanyk 2017; Kong et al. 2018; Zhao, Harrison, and Yu 2024). Traditional approaches, such as API calls, deep links, and Robotic Process Automation (RPA) scripts, offer direct execution paths but suffer from significant limitations, including invalidation from app updates, and a sensitivity to UI changes that hinders their ability to adapt (Bridle and McCreath 2006; Guerreiro, Gamboa, and Jorge 2008; Kennedy and Everett 2011; Le et al. 2020).
The rapid development of LLM-based GUI agents significantly overcomes traditional automation challenges by simulating human interaction through visual or multimodal inputs (Gou et al. 2024; Cheng et al. 2024; Baechler et al. 2024; Liu et al. 2025a). However, complete reliance on LLMs for fine-grained GUI interaction presents new problems: step-by-step action for routine tasks leads to inefficiency and increased costs; in complex operations, LLM hallucinations or misinterpretations can cause cumulative errors, affecting task success rates and reliability (Wen et al. 2024; Wang et al. 2024a). Zhang et al. (Zhang et al. 2025a) provide a detailed analysis of API and GUI agents; the former gain attention for their stability and programming integration capabilities, while the latter possesses strong adaptability. These considerations regarding the efficiency and robustness of GUI agents lead to an exploration of how to combine LLM with efficient shortcut methods, thereby promoting the rise of research in GUI-Shortcut agents.
GUI-Shortcut Hybrid Operations.
Researchers explore GUI-Shortcut methods to enhance agent capabilities and address the persistent challenges in task completion rate, costs, and operational efficiency for LLM-based GUI agents. For instance, UFO2 (Zhang et al. 2025b), a desktop AgentOS, demonstrates robust task execution through its multi-agent architecture and a unified GUI-API action layer that deeply integrates with operating system and application features. Inspired by such work, AppAgentX (Jiang et al. 2025) improves efficiency by evolving high-level actions into shortcuts based on task execution history. MobileAgent-E (Wang et al. 2025a) uses its self-evolving module, learns and stores general-purpose Tips and reusable Shortcuts from past experiences to continuously improve performance and efficiency. While these efforts focused on the mobile domain are promising, they also highlight the urgent need for a systematic method to evaluate the effectiveness of these emerging hybrid operations on mobile devices.
Benchmark # Inst. DS DL EC SG
AndroidArena 221
AndroidWorld 116
LlamaTouch 495
AndroidLab 138
SPA-Bench 340
MAS-Bench (Ours) 139
Table 1: Comparison of MAS-Bench and other smartphone agent benchmarks with dynamic environments. Column definitions: # Inst. (number of instructions), # DS (support diverse shortcuts), # DL (difficulty level), # EC (efficiency&cost metrics), # SG (evaluate shortcut generation).
Mobile GUI Agent Benchmark.
As agents evolve from GUI-based interaction to hybrid GUI-Shortcut operations, they place new demands on benchmarking. Among existing benchmarks, MobileAgentBench (Wang et al. 2024c) focuses on the foundational GUI navigation and task completion capabilities of LLM-powered mobile agents. Mobile-Eval-E (Wang et al. 2025a) includes complex, real-world mobile tasks that require long-horizon, cross-app interactions. While these benchmarks advance research in various dimensions, there remains a need for a benchmark that can comprehensively evaluate the ability of agents to intelligently discover, decide upon, and execute diverse shortcuts within GUI interactions and also measure the overall task effectiveness of this GUI-Shortcut hybrid operation (Zhang et al. 2025b).
3 MAS-Bench
3.1 Online Evaluation Environment
MAS-Bench is built on a dynamic Android platform to evaluate mobile agent performance on complex tasks comprehensively. Unlike static benchmarks, which are often limited to fixed datasets, our platform enables the real-time assessment of agents’ capacity to select and utilize shortcuts intelligently. This dynamic approach moves beyond simply verifying task outcomes, focusing on the agent’s decision-making process in live scenarios. To ensure real-world relevance, the benchmark incorporates a suite of widely used Android applications covering diverse daily-life scenarios. All evaluations are conducted on a standardized emulator to ensure fairness, eliminating confounding factors such as device variations and environmental inconsistencies.
3.2 GUI-Shortcut Hybrid Action Space
To complete tasks, agents in MAS-Bench operate within a hybrid action space that combines conventional GUI interactions with programmatic shortcuts. This space comprises two primary action types:
GUI Action.
GUI actions simulate direct human interaction with an application’s interface. This action space consists of basic operations such as click, swipe, and type, enabling the agent to navigate and interact with UI elements. Detailed definitions for each action are provided in Appendix A.2. GUI actions are essential for dynamic scenarios with unavailable predefined shortcuts.
Shortcut Action.
Shortcut actions can trigger specific functions or navigate to designated pages, bypassing multi-step GUI actions. In MAS-Bench, shortcuts are categorized into two types: Predefined shortcuts and Agent-generated shortcuts.
As illustrated in Fig. 2, the predefined shortcuts in MAS-Bench primarily consist of APIs, deep-links, and RPA scripts.
  • API: Direct programmatic control that executes specific functions by calling an application’s internal services without UI interaction. This method can alter application state or retrieve data directly, as shown in Fig. 2(a).
  • Deep Link: A specialized URI that targets a specific page or function within an application. As shown in Fig. 2(b), this enables direct navigation, bypassing multi-step GUI actions.
  • RPA Script: An Automation script designed to handle a specific, highly repetitive subtask. Unlike a single action, an RPA script encapsulates a complex workflow of GUI actions, API calls, or deep-links, consolidating a multi-step process into a single, efficient shortcut action, as shown in Fig. 2(c).
The agent-generated shortcuts are created dynamically by the agent itself. They are typically formed by identifying and abstracting repetitive subtasks from the agent’s execution history into new, executable routines. This capability allows the agent to learn from experience and progressively optimize its performance by creating customized shortcuts.
Figure 2: Functional Comparison of APIs, Deep-Links, and RPA Scripts. The figure uses the Amazon app as an example: (a) the open_cart() API call directly opens the shopping cart; (b) the search_product() deep-link directly performs a product search; and (c) an RPA script combines APIs, deep links, and GUI operations to automate a complete workflow.
3.3 Shortcut Knowledge Base
Predefined Shortcuts Knowledge Base.
To validate the feasibility of hybrid GUI-shortcut operations on mobile devices, we construct a predefined knowledge base of commonly used mobile shortcuts. Based on 11 popular apps, it includes 11 APIs, 70 Deep Links, and 7 custom RPA scripts, for a total of 88 predefined shortcuts. The knowledge base provides agents with a rich, queryable resource to select appropriate actions based on task instructions. To ensure real-world relevance, APIs were sourced from official documentation and static analysis of application packages. At the same time, Deep Links were identified by analyzing each application’s declared URI schemes and entry points. The custom RPA scripts were designed to encapsulate common, high-repetition sub-tasks, testing an agent’s ability to leverage complex, pre-built automation routines. More details on the collection and design methodology are available in the Appendix A.3.
Agent-Generated Shortcuts.
Besides predefined shortcuts, MAS-Bench supports evaluating agent-generated shortcuts to test an agent’s learning and abstraction capabilities. In order to demonstrate the effectiveness of MAS-Bench in evaluating agents’ capabilities of generating shortcuts, we have incorporated several contrasting shortcut generation methods for evaluation. We include Macro-level action trajectory replay (abbreviated as action replay) shortcuts that record and replay entire task or sub-task execution trajectories. We also incorporate dynamic shortcuts that require the agent to perform real-time grounding based on execution history, as well as shortcuts generated by the MobileAgent-E framework. This diverse set of generation strategies allows for a comprehensive assessment of an agent’s ability to autonomously create efficient workflows.
3.4 Tasks Design for GUI-Shortcut Hybrid Agent
MAS-Bench comprises 139 complex tasks derived from real-world scenarios across 11 Android applications, each featuring one or more invocable predefined shortcuts. These tasks reflect authentic user needs, with an average of 9.27 steps for single-app and 17.66 for cross-app workflows based on human operation. A detailed list of the tasks is provided in Appendix A.4.
Specifically, tasks in MAS-Bench maintain high real-world relevance. They are designed by considering user needs from daily life, such as searching for products in a shopping app or navigating to a location. The tasks typically involve multiple steps and sub-goals, and usable shortcuts exist within our established knowledge base for each task. However, our shortcut knowledge base for these complex tasks is intentionally incomplete. It offers only basic shortcuts, which are insufficient to complete an entire task independently. This design compels agents to synthesize solutions by combining shortcut invocations with multi-step GUI operations, thus providing a more realistic test of their planning and reasoning abilities.
Figure 3: Evaluation Workflow for Agents’ Shortcut Generation Capability. The process consists of two stages: (a) Shortcut Generation Stage, where the agent creates its shortcut knowledge base; and (b) Quality Evaluation Stage. The Agent-generated shortcuts is imported into a baseline agent for performance testing. The quality of the Agent-generated shortcuts is then measured by comparing this performance against the GUI-only baseline agent and the baseline agent with predefined shortcuts.
To test the agent’s learning and generation capabilities, we introduce scenarios with repetitive sub-tasks. In these scenarios, the knowledge base intentionally lacks predefined shortcuts for these recurrent operations. This design aims to test whether an agent, after completing a sub-task via GUI operations for the first time, can identify the repetitive pattern and then autonomously generate an efficient shortcut for subsequent use.
Figure 4: The pipeline of MAS-Bench. The GUI-Shortcut agent first filters products using the search_product shortcut, selects an item via GUI operations, and then adds it to the cart using the add_to_cart shortcut. The entire process is monitored by an automated evaluation module, which outputs metrics such as success rate and efficiency.
3.5 Systematic Hybrid Agent Evaluation
GUI-Shortcut Hybrid Operation Evaluation.
MAS-Bench is designed to evaluate an agent’s ability to autonomously discover and utilize shortcuts to improve task success, efficiency, and cost-effectiveness. This evaluation mandates that an agent possess sophisticated recognition and decision-making capabilities. Agents must not only identify available shortcuts based on the current task and environment but also critically assess their applicability and efficiency. Furthermore, the agent must make optimal decisions between GUI operations and shortcuts to minimize operational steps and time.
Shortcut Generation Evaluation.
We expect GUI agents not only to rely on predefined shortcuts within the knowledge base but also to possess the capability for shortcut generation. The quality of agent-generated shortcuts serves as a direct indicator of their generative capability. However, a systematic framework to develop and evaluate the effectiveness of these generation methods is currently lacking.
Therefore, MAS-Bench encourages and evaluates the agent’s ability to generate new shortcuts autonomously. As shown in Fig. 3, we design the following evaluation method. First, a self-generation shortcut agent explores the MAS-Bench environment and creates its shortcut knowledge base. To ensure fairness and eliminate interference from the intrinsic abilities of the agent under test, we use a unified baseline GUI agent as the evaluated agent and import these shortcut knowledge bases into the baseline agent for task execution. The baseline agent’s performance on the tasks directly reflects the quality of the imported knowledge base, thus enabling an evaluation of each agent’s shortcut generation capabilities.
3.6 Evaluation Metrics
We use the following key evaluation metrics to measure an agent’s performance on MAS-Bench comprehensively. We encourage agents to efficiently complete tasks by utilizing existing and self-generated shortcuts to achieve lower costs and greater efficiency.
Success Rate (SR).
We use SR to measure whether an agent completes a task. Success is typically defined as the agent reaching the final goal state of the task while satisfying all of its requirements.
Efficiency.
We use the following metrics to evaluate the impact of shortcut actions on agents’ efficiency, measured by the time and operations required to complete a task:
  • Mean Steps (MS): The average number of steps an agent takes to complete a task.
  • Mean Step Ratio (MSR): The ratio of the number of steps an agent takes to complete a task to the optimal number of steps required to complete the task. A lower ratio indicates better efficiency in task execution.
  • Mean Execution Time (MET): The average time an agent takes to complete a task.
Cost and Resource Utilization.
Cost metrics evaluate the computational resources and overhead an agent consumes to complete a task. A high-performing agent should minimize the overall cost by intelligently utilizing or autonomously generating shortcuts. We use the following metrics to evaluate the cost of an agent:
  • Mean kTokens Cost (MToC): The average number of kTokens the LLM consumes during the task execution.
  • Mean Shortcut Call Count (MSC): The average number of shortcuts the agent calls during the task execution.
  • Shortcut Success Rate (SSR): The ratio of successfully executed shortcut calls to the total number of shortcut calls made by the agent.
  • GUI to Shortcut Action Ratio (GSAR): This metric reflects an agent’s operational strategy preference by calculating the ratio of GUI to shortcut operations. A high ratio, when combined with low token consumption and short execution time, provides strong evidence that the agent utilizes shortcuts both efficiently and intelligently.
For automated evaluation, we adapt the methodology from SPA-Bench (Chen et al. 2024), embedding shortcut action descriptions to determine task success or failure. MAS-Bench provides a fair, reproducible, and in-depth assessment of a mobile agent’s comprehensive capabilities in complex GUI-Shortcut tasks through these multi-dimensional evaluation metrics. It specializes in agents’ potential for dynamic adaptation and innovative shortcut generation. By doing so, MAS-Bench offers valuable insights and direction for the future development of mobile agents.
4 MAS: Shortcut-Augmented Hybrid Agents
To evaluate the effectiveness of GUI-Shortcut hybrid operation, we introduce MAS-MobileAgent, which serves as a reference for GUI-shortcut Hybrid Agents. MAS-MobileAgent is based on MobileAgent-V2 (Wang et al. 2024b), which primarily relies on visual screenshots for UI understanding and automation. MAS-MobileAgent retrieves relevant shortcuts from a knowledge base and inputs them, along with perceptual information from screenshots, into the model to inform action decisions.
To test the generalization of our shortcut knowledge base and injection method further, we introduce MAS-T3A. Based on T3A (Rawles et al. 2024), this agent relies on a structured UI tree instead of visual information. By applying the same shortcut knowledge base to this agent with a different perceptual foundation, we aim to demonstrate the versatility of our approach, showing it can be seamlessly integrated into different agent architectures.
5 Experiments
Agent Input
[…]
Efficiency Cost
[…]
$S$
VH
[…]
Single-app Tasks (92 Tasks)
Human - 1.000 - - - -
T3A (Rawles et al. 2024) 0.511 1.056 137.641 346.382 0 0
M3A (Rawles et al. 2024) 0.565 1.064 192.775 155.281 0 0
MobileAgentV2 (Wang et al. 2024b) 0.446 1.058 1013.386 120.212 0 0
MobileAgent-E (Wang et al. 2025a) 0.359 0.818 459.574 88.772 0.378 0.081
MAS-T3A (Ours) 0.576 0.915 129.279 291.391 1.043 0.117
MAS-MobileAgent (Ours) 0.641 0.613 682.547 99.780 1.348 0.345
Cross-app Tasks (47 Tasks)
Human - 1.000 - - - -
T3A (Rawles et al. 2024) 0.340 1.087 257.122 625.970 0 0
M3A (Rawles et al. 2024) 0.383 1.262 411.145 288.833 0 0
MobileAgentV2 (Wang et al. 2024b) 0.170 1.247 2053.133 227.128 0 0
MobileAgent-E (Wang et al. 2025a) 0.064 0.934 469.109 85.859 2.250 0.177
MAS-T3A (Ours) 0.511 0.643 185.911 440.222 2.213 0.185
MAS-MobileAgent (Ours) 0.617 0.829 1441.586 189.836 3.128 0.320
Table 2: Performance comparison of our MAS agents and baseline methods on MAS-Bench with a predefined shortcuts knowledge base. All agents utilize the Gemini-2.5-Pro. $S$ and VH refer to the Screenshot and View Hierarchy (UI Tree) input modalities. MSRS is the Mean Step Ratio on Successful tasks, MET is the Mean Execution Time in seconds, and MToC is the Mean Token Cost in thousands (kTokens). The SSR for the predefined shortcuts is 1.0.
We conduct the experiments in the online, dynamic environment of MAS-Bench and evaluate the performance of GUI-shortcut hybrid mobile agent frameworks. We also employ an evaluation framework to measure the efficiency of agent-generated and hand-crafted predefined shortcuts.
5.1 Experiment Setup
Evaluation with Predefined Shortcut.
We evaluate the performance of GUI frameworks on the 139 tasks in MAS-Bench: 1) Frameworks with GUI-only action, including T3A (Rawles et al. 2024), M3A (Rawles et al. 2024), and MobileAgentV2 (Wang et al. 2024b). 2) self-generated shortcut frameworks, including MobileAgent-E (Wang et al. 2025a). 3) shortcut-augmented Hybrid GUI Agents (GUI-shortcut agents), including MAS-T3A and MAS-MobileAgent mentioned in Sec . 4 with predefined shortcut. We employ Gemini-2.5-Pro as the base model for all frameworks. The evaluation utilizes metrics from Sec . 3.6, with detailed configurations in Appendix A.
Evaluation of Shortcut Generation.
To assess the quality of shortcut knowledge bases generated by different strategies, we conduct a two-stage evaluation on a randomly selected test subset comprising 50% of the tasks in MAS-Bench (see Appendix C.2 for the task list). In the Shortcut Generation Stage, we construct several types of agent-generated shortcuts to form distinct knowledge bases. These are created using two methods: the first leverages MobileAgent-E to generate shortcuts ($S_{\text{MobileAgent-E}}$). The second, based on the historical execution trajectories of M3A, generates three types of shortcuts: task-level action replays ($S_{\text{Replay-Task}}$), subtask-level action replays ($S_{\text{Replay-Subtask}}$), and dynamic shortcuts ($S_{\text{Dynamic}}$ ) that require the model to perform real-time grounding, the construction of the agent-generated shortcut knowledge base is detailed in Sec.  3.3.
In the Quality Evaluation Stage, each knowledge base is mapped to a standardized action space and then integrated into the T3A baseline agent. This method treats the baseline agent’s performance as a direct reflection of the imported knowledge base’s quality, enabling an unbiased evaluation of the different generation strategies.
5.2 Effectiveness of GUI-Shortcut Operation
Based on our experimental results in Table 2, we identify three key findings:
Finding 1: Integrating a predefined shortcut knowledge base significantly enhances agent performance, thereby validating the effectiveness of the hybrid GUI-shortcut operation.
Experiments demonstrate that using just the predefined knowledge base of 88 shortcuts can significantly improve an agent’s task success rate and efficiency. For single-app tasks, compared to the baseline T3A, MAS-T3A improves the success rate (SR) from 51.1% to 57.6% and reduces the token cost from 346.382k to 291.391k. For agents that may terminate prematurely on long-horizon tasks due to errors, such as MobileAgent-E in the Table 2, the Mean Step Ratio on Successful tasks (MSRS) provides a more accurate assessment of the efficiency gains from shortcuts. MAS-MobileAgent achieves an MSRS of 0.613 on single-app tasks, which is significantly lower than the baseline version’s 1.058, indicating that its execution path is closer to the optimal solution. The SR and MS results for Different-level Tasks can be found in the Appendix B.2.
Finding 2: The effectiveness of predefined shortcuts is framework-agnostic and can benefit agents with different input modalities.
Experiments show that MAS-T3A and MAS-MobileAgent benefit from predefined shortcuts, regardless of whether they are based on UI Tree or screenshots. The reason lies in that the use of API, Deeplink, and RPA scripts is independent of the agent’s perception modality. These scripts directly send commands to the operating system or application, bypassing the need for the agent to parse the UI layout in real-time.
Figure 5: Performance comparison of MAS-MobileAgent with and without shortcuts. The base models are Gemini-2.5-Pro and Gemini-2.0-Flash. Data points show the relationship between SR and MET for single-app and cross-app tasks, with circle size representing mean cost. Results demonstrate that shortcuts benefit both models, with more significant improvements for the weaker Gemini-2.0-Flash.
Finding 3: The performance gain of shortcuts is more significant when the base model is weak.
For less capable models like Gemini-2.0-Flash, which often struggle with complex GUI navigation, shortcuts provide a crucial way to bypass failure-prone steps. This directly translates into a substantial increase in task success. For instance, on CROSS-app tasks, augmenting the Gemini-2.0-Flash agent with shortcuts boosts its success rate from 0% to 23.4% (see the Appendix B.3 for detailed experimental results). Shortcuts mainly contribute to efficiency improvements for strong models (e.g., Gemini-2.5-Pro). As shown in Fig. 5, the performance gain of using shortcuts is significantly greater for Gemini-2.0-Flash than for Gemini-2.5-Pro.
5.3 Result of Shortcut Generation
Our experiments include a baseline T3A framework without shortcuts and five shortcut variants: Predefined ($S_{\text{Predefined}}$), Replay-Task ($S_{\text{Replay-Task}}$), Replay-Subtask ($S_{\text{Replay-Subtask}}$), Dynamic ($S_{\text{Dynamic}}$), and MobileAgent-E ($S_{\text{MobileAgent-E}}$). Results demonstrate that shortcut quality significantly impacts both execution accuracy and efficiency.
In our experiments, Predefined shortcuts achieve the best performance, improving the success rate (SR) by 9% over the baseline while maintaining a perfect shortcut success rate (SSR) of 100%. Notably, agents using Predefined shortcuts reduce average execution steps by 25% and decrease total execution time by approximately 16%. Despite having the second-highest shortcut call count per task (1.45), Predefined shortcuts maintain excellent robustness, executing successfully across diverse task configurations. The increased success rate and reduced execution time validate that well-designed shortcuts will enhance GUI task efficiency.
In contrast, Replay-Task shortcuts exhibit poor robustness with only a 10% success rate, indicating high susceptibility to environmental variations. This fragility negatively impacts both task completion rates and execution efficiency, underscoring that while effective shortcuts boost GUI performance, poorly designed ones are counterproductive.
Our findings reveal substantial room for improvement in model-generated shortcuts. Among all variants, only Predefined and MobileAgent-E shortcuts improve model accuracy, with only Predefined shortcuts simultaneously enhancing both execution speed and success rate. These results indicate current limitations in shortcut generation capabilities within our framework. Future research should focus on developing more efficient, robust shortcuts with higher utilization rates that effectively reduce execution steps.
Shortcut
[…]
Human - - 1.00 - -
Baseline 0.43 - 0.96 - 188.93
$S_{\text{Predefined}}$ 0.52 1.00 0.71 1.45 152.15
$S_{\text{Replay-Task}}$ 0.34 0.10 0.91 3.04 244.61
$S_{\text{Replay-Subtask}}$ 0.43 0.73 1.13 1.22 236.67
$S_{\text{Dynamic}}$ 0.38 0.75 0.82 0.91 216.24
$S_{\text{MobileAgent-E}}$ 0.49 0.71 1.00 1.01 224.87
Table 3: The results of different shortcut generation methods. Column definitions: # SR (success rate), # MSRS (Mean Step Ratio on Successful tasks), # MSC (Mean Shortcut Call Count), # SSR (Shortcut Success Rate), # MET (Mean Execution Time).
6 Conclusion
In this paper, we introduce MAS-Bench, the first benchmark designed to evaluate the effectiveness of GUI-shortcut hybrid mobile agents. Furthermore, MAS-Bench provides a framework to access the quality and validity of shortcuts that are autonomously generated by the agent. Our experiments demonstrate that the GUI-shortcut hybrid operational approach enhances both the success rate and efficiency of task execution. Moreover, we also validate that our benchmark effectively measures the quality and robustness of the agent-generated shortcuts.
Limitations and Future Work. Despite our efforts to provide a comprehensive evaluation, the range of comparative baselines remains limited due to the currently underexplored research in Mobile GUI-shortcut hybrid Agents. Therefore, we hope to pay more attention to the efficiency of mobile GUI agents in the future, promoting research into GUI-shortcut Hybrid Mobile Agents.
7 Acknowledgments
This work was supported by Huzhou Natural Science Foundation under Grant 2024YZ01 and in part by the Postdoctoral Fellowship Program of CPSF under Grant GZC20241491.
Appendix A MAS-Bench Environment
A.1 Observation Space
MAS-Bench evaluates mobile agents within a highly standardized and controllable environment using an Android Virtual Device (AVD). Specifically, the environment is based on a Pixel 7 emulator running Android 15 (API Level 35). Its screen resolution is set to 1080 x 2400 pixels, and the screen density is 420 dpi. The emulator has 24GB of RAM and 24GB of storage. The environment can support parallel execution of multi-threaded evaluations.
A.2 Details of Action Space
Action Definition
Tap ($x$, $y$) Tap at coordinates ($x$, $y$).
Type (text) Enter text content into an input field.
Swipe ($x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$) Swipe from ($x_{1}$, $y_{1}$) to ($x_{2}$, $y_{2}$).
Home Go to the home screen.
Back Go back to the previous app screen.
Stop Complete the current task.
Table 4: Examples of the basic action space in MAS-Bench.
The MAS-Bench environment facilitates agent-device interaction through a low-level interface based on the Android Debug Bridge (ADB). Crucially, the specific action space available to an agent during evaluation ultimately depends on its architecture and capabilities. Table 4 shows examples of the basic action set.
A.3 Details of Predefined Shortcuts Knowledge Base
Details of Predefined Shortcuts.
Our shortcut knowledge base was created to provide agents with an authentic yet challenging environment. We collect and design shortcuts for 11 selected applications, ensuring their relevance to real-world usage patterns.
  • APIs Collection: The API shortcuts in our knowledge base are application-specific internal APIs. We identify these APIs through official Android documentation. We employ static analysis of the application packages for internal APIs that lack public documentation to identify key functional endpoints.
  • Deep-Links Collection: We collect deep-links by analyzing each application’s AndroidManifest.xml file. This file typically declares the URL schemes and paths the application can handle, providing direct mappings to specific pages or functions.
  • RPAs Script Design: RPA scripts are designed for common, highly repetitive sub-tasks (e.g., adding an item with specific options to a shopping cart). These scripts can be executed as a single action, serving as a high-level shortcut. This design enables us to test an agent’s ability to utilize complex, pre-built automation routines.
Details of Agent-Generated Shortcuts.
We incorporate several methods to create agent-generated shortcuts. The primary purpose for including this variety of shortcut types is to validate the effectiveness of the MAS-Bench framework itself in evaluating an agent’s shortcut generation capabilities. By creating test cases with diverse characteristics, we can verify our benchmark’s ability to assess different shortcut generation strategies.
  • Macro-level action trajectory replay (action replay) Shortcuts: This category of shortcuts is constructed by recording and replaying an agent’s historical action trajectories. They execute a fixed sequence of low-level actions, such as clicks and swipes at specific coordinates, to automate a previously completed workflow. This method is implemented at two distinct granularities:
    • Task-Level Replay: Captures and replays the entire action sequence of a full task, from start to finish. This results in a single, high-level macro that is highly specific to one complete workflow.
    • Subtask-Level Replay: Identifies and abstracts recurrent, multi-step operations within a larger task. This approach generates shorter, more modular shortcuts that automate common sub-workflows.
  • Dynamic Shortcuts: Unlike static macro-replays, these shortcuts are adaptive and robust to UI changes. Instead of replaying hardcoded coordinates, they utilize real-time grounding to semantically identify and interact with target UI elements based on the current screen context. This allows them to function correctly even if the position or layout of UI elements changes.
Ultimately, this diverse suite of shortcut types validates that MAS-Bench provides a comprehensive framework for evaluating agents’ shortcut generation capabilities.
Agent Base Model
[…]
Efficiency Cost
[…]
Single-app Tasks (92 Tasks)
MobileAgentV2 Gemini-2.0-Flash 0.250 1.280 472.426 40.406 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.402 0.719 335.181 39.597 3.565 0.505
MobileAgentV2 Gemini-2.5-Pro 0.446 1.058 1013.386 120.212 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.641 0.613 682.547 99.780 1.348 0.345
Cross-app Tasks (47 Tasks)
MobileAgentV2 Gemini-2.0-Flash 0 - 952.660 106.501 0 0
MAS-MobileAgent Gemini-2.0-Flash 0.234 0.738 403.798 71.350 5.362 0.709
MobileAgentV2 Gemini-2.5-Pro 0.170 1.247 2053.133 227.128 0 0
MAS-MobileAgent Gemini-2.5-Pro 0.617 0.829 1441.586 189.836 3.128 0.320
Table 5: Evaluation results of different base models. MSRS is the Mean Step Ratio on Successful tasks, MET is the Mean Execution Time in seconds, and MToC is the Mean Token Cost in thousands (kTokens).
A.4 Details of Tasks in MAS-Bench
We have developed a specialized benchmark suite to validate the feasibility of GUI-shortcut hybrid actions for mobile agents and establish a framework for the comprehensive evaluation of an agent’s ability to utilize and generate shortcuts.
This suite comprises 139 complex tasks strategically designed across 11 popular real-world applications. The tasks are categorized into 92 single-app tasks and 47 cross-app workflows, with their difficulty distribution illustrated in Table 6. We have also curated a corresponding knowledge base of predefined shortcuts to support these tasks. Table 9 presents a breakdown of the selected applications and the number of single-app tasks and shortcuts associated with each.
Difficulty Level Proportion (%) Human Steps
Level 1 27.3 6.2
Level 2 47.5 12.6
Level 3 25.2 17.6
Table 6: Distribution of task difficulty levels in MAS-Bench. The table shows the proportion of tasks at each level and the average number of steps required for a human to complete them.
Appendix B Details of Predefined Shortcut Evaluation
B.1 Details of MAS-MobileAgent and MAS-T3A
Compared to MobileAgentV2, MAS-MobileAgent acquires the standard screenshot and UI elements and retrieves relevant and potentially applicable shortcuts from the shortcut knowledge base, based on the current task goal and interface context. This retrieved shortcut information is injected as an additional context into the prompt provided to the base model, along with screen information and the task goal.
MAS-T3A operates on a similar principle, but its core distinction lies in its input modality: it relies on the structured UI tree, whereas MAS-MobileAgent processes visual screenshots. By evaluating both agents, we aim to demonstrate that the effectiveness of predefined shortcuts is framework-agnostic and can benefit agents with fundamentally different perceptual modalities.
B.2 Success Rate on Different-Level Tasks
As shown in Table 7, an analysis of metrics such as MSRS reveals that the GUI-shortcut hybrid agent, by utilizing the predefined shortcuts knowledge base, demonstrates improvements in both success rate and efficiency across tasks of varying difficulty levels when compared to the GUI-only agent.
Agent DL SR MS MSRS
T3A Level 1 71.1 7.5 0.93
Level 2 34.8 17.6 1.04
Level 3 42.9 21.2 1.07
M3A Level 1 65.8 7.7 1.02
Level 2 51.5 17.0 1.19
Level 3 31.4 25.3 1.04
MobileAgentV2 Level 1 52.6 18.2 1.04
Level 2 34.8 30.6 1.10
Level 3 17.1 46.8 1.24
MobileAgentE Level 1 44.7 8.8 0.84
Level 2 20.0 5.5 0.82
Level 3 17.1 6.2 0.60
MAS-T3A Level 1 76.3 6.2 0.78
Level 2 53.0 13.2 0.82
Level 3 37.1 20.7 0.96
MAS-MobileAgent Level 1 84.2 8.0 0.54
Level 2 65.2 18.8 0.76
Level 3 37.1 32.2 0.79
Table 7: Performance comparison of our MAS agents and baseline methods on MAS-Bench across different difficulty levels. Column definitions: # DL (difficulty level), # SR (success rate), # MS (Mean step), # MSR (Mean Step Ratio on Successful tasks).
B.3 Influence of Base Model Capability
To assess the impact of GUI-shortcut hybrid actions on base models with varying capabilities, we conducted experiments using Gemini 2.0-Flash and Gemini 2.5-Pro. Our experiments, detailed in Table 5, reveal that the performance gain from shortcuts is significantly more pronounced when the base model is weak. For the less capable Gemini 2.0-Flash, shortcuts are critical for task completion. This is most apparent in cross-app tasks, where shortcuts boost the Success Rate (SR) from 0 to 0.234, making previously impossible tasks solvable. Efficiency also improves dramatically, with Mean Execution Time (MET) cut by over 57% in these scenarios.
Appendix C Details of Shortcut Generation Evaluation
C.1 Shortcut Generation Method
To validate the effectiveness of the MAS-Bench framework in evaluating an agent’s shortcut generation capabilities, we generated a diverse set of agent-generated shortcuts using various methods.
We employ M3A as our baseline agent to generate different types of shortcuts. Without altering its core operational logic, M3A autonomously explores the environment and generates task execution trajectories, utilizing Gemini-2.5-Pro as its base model. Upon completing each task, we construct a prompt containing the task instruction, the whole action history, and the reasoning for each step. This prompt is then provided to Gemini-2.5-Pro, generating the corresponding shortcuts.
Examples of the resulting shortcut types are shown in Fig. 6. As detailed in Table 8, this process resulted in 46 Subtask-level Macro-Replay, 45 Dynamic, 39 Task-level Macro-Replay, and 36 MobileAgent-E shortcuts.
Shortcut Type Number of Shortcuts
$S_{\text{Replay-Task}}$ 39
$S_{\text{Replay-Subtask}}$ 46
$S_{\text{Dynamic}}$ 45
$S_{\text{MobileAgent-E}}$ 36
Table 8: Number of agent-generated shortcuts by different methods.
Figure 6: Examples of the resulting shortcut types. Action Replay shortcuts (Task-Level and Subtask-Level) use a sequence of actions with fixed indices, while Dynamic Shortcuts use variable arguments that correspond to UI elements.
C.2 Subset Task
To validate our framework for evaluating an agent’s shortcut generation capability, we randomly selected a subset of 50% from the 139 tasks in MAS-Bench. This subset comprises a total of 69 tasks, including 46 single-app and 23 cross-app tasks.
Appendix D Bad Case Analysis
We further analyze the performance of the GUI-shortcut hybrid agent on failed tasks, with a particular focus on cases where the GUI-only agent succeeds, but the GUI-shortcut hybrid agent fails. Our analysis of these failure cases reveals three primary
Figure 7: Examples of GUI-shortcut hybrid agent failure cases. (a) Selection and Planning Error: The agent incorrectly invokes search_hotel() shortcut instead of searching for attractions, demonstrating failure in mapping natural language instructions to appropriate shortcuts. (b) Behavioral and Adaptation Error: The agent repeatedly calls open_shorts() shortcut instead of switching to a GUI-only action to select the video, showing an inability to adapt operational mode based on task state.
Selection and Planning Errors.
This category of error pertains to flaws in the agent’s task comprehension and strategic planning. These errors typically occur when the agent fails to accurately map the natural language instruction to the correct shortcut, resulting in the invocation of an irrelevant shortcut that leads to task failure.
For instance, in an experiment with MAS-MobileAgent using Gemini-2.0-Flash (Fig. 7(a)), the agent was given the instruction: “Navigate to the Attractions section. Search for attractions near Berlin for the 6th of next month.” After launching the application, the agent made a planning error by incorrectly invoking the search_hotel() shortcut, which was irrelevant to the “Attractions” goal, thus causing the task to fail.
Behavioral and Adaptation Errors.
This error class occurs when the agent fails to evaluate the outcome of an action and adjust its subsequent action. The deficiency typically stems from an inability to learn from the action histories and modify future behavior accordingly.
As shown in Fig. 7(b), when tasked to “Get the search results for the short videos Stephen Curry. Select the first Short in the results and like it. Leave a comment good shot!”, The agent successfully uses shortcuts to search for the video. However, at the step requiring it to select the first video, it fails to switch to a GUI-based action and instead erroneously calls the open_shorts() shortcut again. This results in an unproductive loop, demonstrating a failure to adapt its operational mode based on the task’s state.
Execution and Formulation Errors.
This type of error is specific to the invocation of agent-generated shortcuts. As noted in the main text, the execution success rate for these shortcuts is considerably low. Failures typically result from the shortcut being poorly formulated during the generation phase, rendering it inherently flawed or not robust enough to handle slight variations in the UI during execution.
App name Tasks Shortcuts Shortcut Example
YouTube 11 5 youtube.search_video_query("cat")
youtube.open_shorts()
Amazon 10 10 amazon.open_amazon()
amazon.open_cart()
Booking.com 10 10 booking.open_hotelid()
booking.open_my_trips()
Google Calendar 9 2 calendar.open_calendar_main()
calendar.create_event(title="Meeting")
Google Maps 9 8 google_map.navigate_to_address("Beijing")
google_map.search_address("Square")
BBC 8 16 bbc.open_news()
bbc.open_live()
Fitbit 8 21 fitbit.open_fitbit_main()
fitbit.open_sleep_log()
Gmail 8 3 gmail.open_gmail()
gmail.open_setting()
Chrome 7 4 chrome.search_query("Python")
chrome.open_incognito()
Yelp 8 6 yelp.open_yelp_main()
yelp.open_events()
Contacts 4 3 contacts.open_contacts_main()
contacts.open_contact_starred()
Table 9: List of MAS-Bench apps and number of single-app tasks and shortcuts for each one.